Use Euclidean algorithm to determine the gcd (4076, 1024)


1
Expert's answer
2022-05-23T16:05:06-0400

1. 4076/1024=3 whith remainder 1004

So 4076=3x1024+1004

2. 1024/1004=1 whith remainder 20

1024=1x1004+20

3. 1004/20=50 whith remainder 4

1004=20x50+4

4. 20/4=5 whith remainder 0.

So gcd of 4076 and 1024 is 4